# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

### v2

@@
Abstract.
Portrait image animation using audio has rapidly advanced, but challenges remain in efficiently fusing multimodal inputs while ensuring temporal and portrait consistency with minimal computational cost. To address this, we present LetsTalk, a LinEar diffusion TranSformer for Talking video synthesis. LetsTalk incorporates a deep compression autoencoder to obtain efficient latent representations, and a spatio-temporal-aware transformer with efficient linear attention to effectively fuse multimodal information and enhance spatio-temporal consistency. We systematically explore and summarize three fusion schemes, ranging from shallow to deep fusion. We thoroughly analyze their characteristics, applicability, and trade-offs, thereby bridging critical gaps in multimodal conditional guidance. Based on modality differences of image, audio, and video generation, we adopt deep (Symbiotic Fusion) for portrait to ensure consistency, and shallow (Direct Fusion) for audio to align animation with speech while preserving motion diversity. To maintain temporal consistency in long-duration video generation, we propose a memory bank mechanism that preserves inter-clip dependencies, effectively preventing degradation across extended sequences. Furthermore, we develop a noise-regularized training strategy that explicitly compensates for DDPM sampling artifacts, significantly improving the model’s robustness in continuous generation scenarios. Our extensive experiments demonstrate that our approach achieves state-of-the-art generation quality, producing temporally coherent and realistic videos with enhanced diversity and liveliness, while maintaining remarkable efficiency through its optimized model design with 8$\times$ fewer parameters.
Talking Head Generation, Multimodal Guidance, Long-duration Generation, Linear Diffusion Transformer, Deep Compression Autoencoder
ccs: Computing methodologies Procedural animationccs: Computing methodologies Computer vision tasksccs: Computing methodologies Neural networks
Figure 1. We introduce LetsTalk, a diffusion-based transformer for audio-driven portrait animation. Given a reference image and audio, LetsTalk generates realistic videos with synchronized mouth motions. As shown in the Left figure, each column corresponds to the same audio, demonstrating consistent and accurate lip movements. The Right figure compares generation quality and inference time on the HDTF dataset, where circle size represents model parameters. LetsTalk achieves superior quality and efficiency, outperforming methods like Hallo and AniPortrait. Notably, our base version (LetsTalk-B) matches Hallo’s performance with 8$\times$ fewer parameters.
@@ Abstract.
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
@@
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc.. Previous works (Zhang et al., 2023a; Wei et al., 2024) use a 3D morphable model (3DMM) (Blanz and Vetter, 1999; Li et al., 2017; Bao et al., 2021) to represent reference portrait and drive them with audio signals. These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair. In contrast, audio-to-2D facial motion methods (Chen et al., 2019; Prajwal et al., 2020; Zhou et al., 2020; Cheng et al., 2022; Zhou et al., 2021) directly align audio with facial dynamics via latent space learning. However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
@@
Recently, diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022) have emerged as a powerful paradigm for generating realistic 2D images (Rombach et al., 2022; Zhang et al., 2023b) and have demonstrated remarkable success in video generation (Ma et al., 2024; Lu et al., 2023). Notably, AnimateAnyone (Hu, 2024) introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion. Building on this, recent methods (Xu et al., 2024b; Chen et al., 2024; Cui et al., 2024) have extended this approach by replacing pose signals with audio signals for talking head generation. These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
@@
Despite these advancements, several critical challenges remain unaddressed. First, current approaches suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific requirements; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity. Second, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio. Lastly, while existing methods attempt to achieve long-duration portrait animation through sequential frame synthesis, they fail to address the progressive accumulation of residual noise in generated frames. This phenomenon inevitably introduces cumulative artifacts (e.g., facial distortions or flickering) that amplify over extended sequences, ultimately degrading output fidelity.
In this work, we introduce LetsTalk, a linear diffusion transformer for efficient generation of realistic, long-duration talking videos under multimodal conditions. Built upon a diffusion transformer as the video generation backbone, LetsTalk replaces conventional text-to-image architectures with an efficient design tailored for audio-driven animation. Key to our approach is a deep compression autoencoder that compresses input frames 32×, drastically reducing the number of latent tokens while preserving critical visual details. Each transformer block incorporates dedicated temporal and spatial modules, where linear attention (Xie et al., 2024a) ensures efficient modeling of intra-frame and cross-frame coherence without sacrificing computational efficiency.
@@
Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio. To better exploit multimodality to guide video generation, we first summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities, and further investigate the underlying factors influencing such adaptability—thereby bridging theoretical gaps in conditional guidance modeling. To better guide video generation effectively, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency due to its visual similarity to the video. For audio, we employ Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion).
To enhance long-duration video generation, we incorporate a memory bank mechanism to maintain contextual continuity across sequential video sequences. Furthermore, we propose a novel noise-regularized training strategy and a patch drop augmentation to align the memory bank between training and inference. By explicitly simulating the residual noise accumulation that occurs during inference (where errors propagate as the video extends), our approach learns to compensate for such artifacts during training for robust long-sequence generation.
We summarize the contributions of this work as follows:
  • We present LetsTalk, a linear diffusion transformer for audio-driven portrait animation, which employs efficient linear attention for spatiotemporal modeling and integrates a deep compression autoencoder to achieve 32 $\times$ latent space reduction while preserving visual fidelity;
  • We systematically analyze three multimodal fusion schemes, characterize their adaptation properties, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality guidance;
  • We enable long-duration video generation via a memory bank framework with noise-regularized training, aligning memory states between training and inference to mitigate cumulative artifacts in extended sequences;
  • Our experiments demonstrate our method’s superiority in generating realistic, temporally coherent videos with enhanced diversity, achieving an optimal balance between computational efficiency and output quality.
2. Related Work
2.1. Diffusion Models for Video Generation
@@
Diffusion models have demonstrated exceptional potential across various domains, particularly in image (Ho et al., 2020; Dhariwal and Nichol, 2021) and video generation (Guo et al., 2023; Hu, 2024; Ho et al., 2022b; Zhu et al., 2024). A notable example is Stable Diffusion (Rombach et al., 2022), which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets (Schuhmann et al., 2022). This approach has been extended to more complex tasks like video generation (Guo et al., 2023; Hu, 2024), where models such as Video Diffusion Models (VDM)(Ho et al., 2022b) apply the diffusion process with space-time factorized training, and ImagenVideo(Ho et al., 2022a) enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer (Wang et al., 2024b), AnimateDiff (Guo et al., 2023), and VideoCrafter (Chen et al., 2023b), focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023), which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte (Ma et al., 2024) and VDT (Lu et al., 2023), have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
2.2. Audio-driven Portrait Image Animation
@@
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert (Prajwal et al., 2020) achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet (Chen et al., 2019) and MakeItTalk (Zhou et al., 2020) introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker (Zhang et al., 2023a), Genface (Ye et al., 2023), and DiffTalk (Shen et al., 2023) further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads (Stypułkowski et al., 2024), DreamTalk (Ma et al., 2023), and VividTalk (Sun et al., 2023), which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 (Xu et al., 2024a), EMO (Tian et al., 2024) and AniPortrait (Wei et al., 2024) have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker (Liu et al., 2024) introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait (Guo et al., 2024) extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo (Xu et al., 2024b; Cui et al., 2024), EchoMimic (Chen et al., 2024) and X-Portrait (Xie et al., 2024b) have adopted the latent diffusion framework pioneered in AnimateAnyone (Hu, 2024), adding more control over elements such as lip movements, expressions, and poses.
2.3. Conditionally Guided Diffusion Model
@@
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms (Ramesh et al., 2022; Avrahami et al., 2022; Brooks et al., 2023; Yang et al., 2024; Khachatryan et al., 2023). ControlNet (Zhang et al., 2023b) enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone (Hu, 2024) introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 (Esser et al., 2023) integrates structural guidance through depth maps and cross-attention, while MCDiff (Chen et al., 2023a), LaMD (Hu et al., 2023), and MotionCtrl (Wang et al., 2024a) focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits (Shen et al., 2023; Stypułkowski et al., 2024). Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information (Tian et al., 2024; Xu et al., 2024b). While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio). To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
@@
Figure 2. The illustration of our method. LetsTalk integrates a deep compression autoencoder that efficiently reduces spatial redundancy while preserving temporal features, and transformer blocks equipped with dual temporal and spatial attention modules, enabling the capture of intra-frame spatial details while establishing temporal correspondences across frames. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding. Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding. To support long-duration video generation, we introduce a memory bank module to preserve temporal consistency across video sequences. Additionally, we devise a novel training strategy to to align memory bank between training and inference. By explicitly simulating the residual noise during inference, our approach learns to compensate for such artifacts, ensuring robust long-duration generation.
3. Method
@@
We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig. 2. In this section, we first detail our efficient transformer architecture in Sec. 3.1. Next, we outline multimodal fusion schemes in Sec. 3.2 and describe our strategy for integrating reference portraits with audio in Sec. 3.3. Finally, we introduce our noise-regularized training approach for long-duration video generation Sec. 3.4.
3.1. Efficient Spatial-Temporal Transformer with Compression Autoencoder
@@
Deep Compression Autoencoder: Building upon latent diffusion models, we employ a deep compression autoencoder (Xie et al., 2024a) that downsamples images by a factor of 32 (vs. 8 in conventional VAEs), while maintaining reconstruction quality comparable to SDXL’s autoencoder through advanced architectural designs. This design is motivated by two key insights: (1) high-resolution video frames contain substantial redundant information that can be safely compressed, and (2) for audio-driven generation, separating compression from denoising enables specialized processing - the autoencoder preserves crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment.
The compressed latent feature $\mathcal{F}\in\mathbb{R}^{F\times H\times W\times C}$ is processed through non-overlapping patches, where spatial-temporal positional embeddings (Vaswani, 2017) preserve structural relationships. This compact representation enables efficient transformer processing while reducing memory requirements by 4$\times$ compared to conventional approaches.
@@
Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal. While the original DiT (Peebles and Xie, 2023) processes individual frames independently, our video generation framework extends this through interleaved spatiotemporal fusion. The naive approach of flattening $\bm{Z}$ into $(F\times P,d)$ tokens with standard attention incurs prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution video generation.
Our architecture employs a dual-block decomposition strategy that alternates between spatial and temporal modeling, significantly reducing computational complexity from $\mathcal{O}(F^{2}P^{2})$ to $\mathcal{O}(F^{2}+P^{2})$. To reduce computational complexity, we replace the traditional softmax attention with ReLU linear attention (Katharopoulos et al., 2020). For spatial modeling, the input is first reshaped into $(F,P,d)$ dimensions to compute intra-frame attention using:
(1) $\bm{Z_{s}}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}$
where $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$, $\phi(x)=\text{ReLU}(x)$. The linear attention mechanism further optimizes this by reusing shared terms across all queries: $\sum_{j=1}^{N}\text{ReLU}(\bm{K}_{j})\bm{V}_{j}^{T}\in\mathbb{R}^{d\times d}$.
For temporal modeling, the tensor to $(P,F,d)$ dimensions. Since $F\ll P$, we preserve the expressive power of standard softmax attention for inter-frame interactions: $\bm{Z_{t}}=\text{Attn}(\bm{Q}_{t},\bm{K}_{t},\bm{V}_{t})=\frac{\bm{V}_{t}(\bm{K}_{t}^{\top}\bm{Q}_{t})}{\sqrt{d}}$. Through this hybrid decomposition with linear attention, the overall complexity is reduced to $\mathcal{O}(F^{2}+Pd))$ achieving near-linear efficiency while preserving global receptive fields.
@@
Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance. Following SANA (Xie et al., 2024a), we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU. The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality. Our experiments demonstrate this architecture achieves superior computational efficiency without compromising visual quality or temporal consistency compared to quadratic-complexity baselines in Fig. 1.

### v3

@@
Abstract
@@
Index Terms:
Talking Head Generation, Long-duration Generation, Multimodal Guidance, Linear Diffusion Transformer, Deep Compression Autoencoder
[…]
I Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc. Despite recent progress, long-duration talking video synthesis still faces persistent challenges in simultaneously achieving high video quality, portrait and temporal consistency, and computational efficiency. As video length increases, issues such as visual degradation, loss of identity consistency, temporal incoherence, and error accumulation become increasingly prominent, severely impacting the realism and reliability of generated results.
@@ I Introduction
Previous works [1, 2] use a 3D morphable model (3DMM) [3, 4, 5] to represent reference portraits and drive them with audio signals. These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair. In contrast, audio-to-2D facial motion methods [6, 7, 8, 9, 10] directly align audio with facial dynamics via latent space learning. However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
@@
Recently, diffusion models [11, 12, 13] have emerged as a powerful paradigm for generating realistic 2D images [13, 14] and have demonstrated remarkable success in video generation [15, 16]. Notably, AnimateAnyone [17] introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion. Building on this, recent methods [18, 19, 20] have extended this approach by replacing pose signals with audio signals for talking head generation. These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
@@
However, several critical challenges remain unaddressed, particularly in the context of long-duration video generation. As the length of generated videos increases, existing methods struggle to maintain high visual quality, portrait consistency, and temporal coherence. Progressive accumulation of residual noise and artifacts (e.g., facial distortions or flickering) becomes increasingly severe over extended sequences, ultimately degrading output fidelity and realism. In addition to these long-duration challenges, current approaches also suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific tasks; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity. Furthermore, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio.
In this work, we introduce LetsTalk, a diffusion transformer framework designed for efficient and high-quality generation of realistic, long-duration talking videos under multimodal conditions. A key component for long video generation in LetsTalk is the memory bank mechanism, which enables each generated clip to access information from previous clips, ensuring smooth transitions and contextual continuity. To address the discrepancy between training and inference, we propose a noise-regularized training strategy that explicitly simulates inference-time noise accumulation, thereby mitigating error accumulation and sampling artifacts as the video extends.
[…]
@@
Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio. To better exploit multimodality for guiding long-duration video generation, we first summarize three fusion schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities. We further investigate the underlying factors influencing such adaptability, thereby bridging theoretical gaps in conditional guidance modeling. To effectively guide long-duration video generation, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency throughout extended sequences due to its visual similarity to the video. For audio, we adopt Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion) over long video durations.
We summarize the contributions of this work as follows
  • We propose LetsTalk, a novel diffusion transformer framework for long-duration talking video generation, which simultaneously achieves high video quality, portrait and temporal consistency, and computational efficiency.
  • We introduce a memory bank mechanism combined with a noise-regularized training strategy, which explicitly maintains contextual continuity and effectively mitigates error accumulation and sampling artifacts during long video generation.
  • We design an efficient spatiotemporal modeling architecture, leveraging a deep compression autoencoder and linear attention, which achieves comparable visual fidelity with only 1/8 the parameters of previous methods.
  • We systematically analyze multimodal fusion schemes and adopt deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio, enabling LetsTalk to achieve state-of-the-art visual quality and temporal coherence in talking video generation.
II Related Work
II-A Diffusion Models for Video Generation
@@
Diffusion models have demonstrated exceptional potential across various domains, particularly in image [11, 12] and video generation [22, 17, 23, 24]. A notable example is Stable Diffusion [13], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [25]. This approach has been extended to more complex tasks like video generation [22, 17], where models such as Video Diffusion Models (VDM)[23] apply the diffusion process with space-time factorized training, and ImagenVideo[26] enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer [27], AnimateDiff [22], and VideoCrafter [28], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [29], which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte [15] and VDT [16], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
II-B Audio-driven Portrait Image Animation
@@
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [7] achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet [6] and MakeItTalk [8] introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker [1], Genface [30], and DiffTalk [31] further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads [32], DreamTalk [33], and VividTalk [34], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 [35], EMO [36] and AniPortrait [2] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker [37] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait [38] extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo [18, 20], EchoMimic [19] and X-Portrait [39] have adopted the latent diffusion framework pioneered in AnimateAnyone [17], adding more control over elements such as lip movements, expressions, and poses.
II-C Conditionally Guided Diffusion Model
@@
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [40, 41, 42, 43, 44]. ControlNet [14] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone [17] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 [45] integrates structural guidance through depth maps and cross-attention, while MCDiff [46], LaMD [47], and MotionCtrl [48] focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [31, 32]. Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [36, 18]. While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio). To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
III Methodology
@@
Figure 2: The illustration of our method. LetsTalk is designed for robust long-duration talking video generation. It integrates a deep compression autoencoder to reduce spatial redundancy while preserving temporal features, and transformer blocks with temporal and spatial attention to capture intra-frame details and long-range temporal correspondences. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for the portrait embedding and Direct Fusion for the audio embedding, providing effective multimodal guidance for long video synthesis. The portrait embedding is repeated along the frame axis to match the noise embedding for consistent conditioning. To support long video generation, we introduce a memory bank module to preserve temporal consistency across extended sequences. We also propose a noise-regularized training strategy to align the memory bank between training and inference, enabling stable and high-fidelity long-duration generation.
@@
We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig. 2. This section first introduces the diffusion formulation and transformer backbone in Section III-A. We then detail our long-range consistency mechanism, including the memory bank and noise-regularized training, in Section III-C. Next, we describe the efficient spatiotemporal transformer backbone in Section III-B. Finally, we analyze multimodal fusion schemes and our guidance strategy in Section III-D and Section III-E.
III-A Preliminaries
$\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) [49] in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
III-B Efficient SpatioTemporal Transformer with Linear Attention
@@
Deep Compression Autoencoder: We build upon latent diffusion models and employ a deep compression autoencoder [21] that downsamples images by a factor of 32, maintaining reconstruction quality comparable to SDXL’s autoencoder. This design leverages the redundancy in high-resolution video frames and decouples compression from denoising, allowing the autoencoder to preserve crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment. Compared to conventional autoencoders, this approach significantly reduces memory and computation requirements while maintaining high fidelity.
[…]
@@
Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal. Unlike the original DiT [29], which processes each frame independently, our framework introduces interleaved spatiotemporal fusion to jointly model spatial and temporal dependencies.
A naive approach would flatten $\bm{Z}$ into $(F\times P,d)$ tokens and apply standard self-attention, resulting in prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution videos. To address this, we adopt a dual-block decomposition strategy that alternates between spatial and temporal attention blocks, significantly reducing the computational cost to $\mathcal{O}(F^{2}+P^{2})$.
For spatial modeling, we reshape the input to $(F,P,d)$ and apply ReLU-based linear attention [51] within each frame:
$\bm{Z}_{s}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}},$ (3)
where $\phi(x)=\text{ReLU}(x)$ and $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$.
For temporal modeling, we transpose the tensor to $(P,F,d)$ and apply standard softmax attention across frames, allowing us to preserve the expressive power of full attention for inter-frame interactions.
By interleaving these spatial and temporal attention blocks, our architecture achieves near-linear efficiency while maintaining a global receptive field, enabling scalable and effective modeling of high-resolution, long-duration talking videos.
@@
Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance. Following SANA [21], we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU. The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality. This enhances local feature capture and compensates for the weaker locality of linear attention.
Adaptive Integration Following VDT [16], we insert an adaptive fully-connected layer between attention layers to fuse temporal dynamics into spatial representations:
$\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$ (4)
Figure 3: The illustration of the long-duration generation.
III-C Memory Bank and Noise-Regularized Training for Long-Term Consistency
In diffusion-based video generation, maintaining temporal alignment and consistency across long sequences is essential. However, as the video extends, the accumulation of residual noise and sampling artifacts can threaten temporal coherence and visual quality. This is further exacerbated by the mismatch between training and inference: during training, the memory bank is constructed from ground-truth frames, while during inference, it is updated with generated frames.
Memory Bank: To address these challenges, we introduce a memory bank that stores information from previously generated clips. Unlike previous frame-by-frame generation approaches [31, 32], our end-to-end diffusion transformer generates an entire video clip in a single sampling process. The memory bank is patchified and temporally transposed, resulting in $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ denotes the number of frames selected from prior clips. The memory bank provides prior information to the current generation through Symbiotic Fusion with $\bm{Z_{t}}$, leveraging the scalability of the Transformer architecture. This design ensures smooth transitions and temporal consistency across video segments, and is especially advantageous for long-duration video generation.
Theoretical Motivation: We observed that DDPM-based sampling cannot fully remove noise due to two main factors:
  • Model prediction error: The predicted noise $\epsilon_{\theta}(x_{t},t)$ only approximates the true noise $\epsilon$, leaving some residual noise after denoising.
  • Sampling stochasticity: At each reverse step, new Gaussian noise is added:
    $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)+\sigma_{t}z,\quad z\sim\mathcal{N}(0,I)$ (5)
    The term $\sigma_{t}z$ introduces inherent randomness.
Consequently, even after finite-step denoising, diffusion models retain residual noise, which accumulates as the video sequence extends. This motivates our noise-regularized memory bank, which differs from previous long video generation methods.
Noise-Regularized Training Strategy. To bridge the gap between training and inference, we propose a noise-regularized training strategy. We perturb the memory bank with controlled Gaussian noise, $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$, simulating inference-time noise accumulation. Specifically, we randomly sample a diffusion timestep $t$ from the first 50 steps and add the corresponding noise to the memory bank during training. Additionally, inspired by Hallo2 [20], we apply patch-drop augmentation by randomly masking 25% of spatial patches in the memory bank, further enhancing model robustness.
This deliberate corruption enables the model to learn to correct for artifacts and maintain temporal dependencies. Our framework thus achieves robust and stable long-duration video synthesis, as demonstrated in Fig. 3.